OpenStudy (anonymous):

For the given function, state the amplitude and the maximum output for the function: ƒ(t) = 2sin(t). A.) amp=2 max=4 B.) amp=4 max=4 C.)amp=4 max=2 d.) amp=2 max=2

OpenStudy (perl):

The general sinusoidal function is y = A * sin( 2pi / m ( x - p.s.) ) + v.s. where A = amplitude p.s. = phase shift v.s. = vertical shift m = period

OpenStudy (anonymous):

Ive never heard of that function

OpenStudy (anonymous):

@zepdrix sorry to bother you again but can you please explain this to me if you know how to do it cause Perl has confused me more. :/ Sorry

zepdrix (zepdrix):

Hmm I'm not sure if I could explain it much better. I think he gave you just a little bit too much information for this problem, so it's a bit of a brain overload hehe.

zepdrix (zepdrix):

When we have this function,\[\Large\rm y=\sin t\]It's a periodic function, it reaches a maximum output of 1, a minimum output of -1, and it repeats it's pattern over and over.

zepdrix (zepdrix):

\[\Large\rm y=A \sin t\]As Perl said, This A represents magnitude. When a number is put in place of this A, it will change the maximum and minimum outputs of the function.

OpenStudy (anonymous):

Oh i think i get it now. Perl just put it in big mathmatical sense.

zepdrix (zepdrix):

He called it amplitude (blah sorry). That makes more sense, since it's the letter A ;)

OpenStudy (anonymous):

True

zepdrix (zepdrix):

What do you think for this one? Are there two options we can cross off immediately since we know the value of the amplitude?

OpenStudy (anonymous):

b and c? right?

zepdrix (zepdrix):

Good good good. So we narrowed it down to A or D.

OpenStudy (anonymous):

Id say a.

OpenStudy (anonymous):

no wait!

zepdrix (zepdrix):

If \(\Large\rm y=\sin t\) has a maximum output of 1, what will the new maximum output be if we multiply that by 2?

OpenStudy (anonymous):

2?

zepdrix (zepdrix):

1*2 = 2. Good good good. :)

OpenStudy (anonymous):

oh okays. Why did this happen to be super easy... -_- oh well Thank you again! Youre very kind :)

zepdrix (zepdrix):

hehe

OpenStudy (perl):

sorry i gave more info than you need. but keep that formula for later, it may come in handy